OpenStudy (anonymous):

solve the equation for x square root of x-3= 4 show work?

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

\[x^{2}+9=16\]

OpenStudy (anonymous):

that doesnt work?

OpenStudy (anonymous):

If u square a square root, you get what is under the radical

OpenStudy (anonymous):

x * sqrt(x-3) - square this

OpenStudy (anonymous):

No, it's just sqrt(x-3)

OpenStudy (anonymous):

sqrt(x-3)=4 i need to find out what x is

OpenStudy (anonymous):

Square it..

OpenStudy (anonymous):

If u square a square root, you get what is under the radical

OpenStudy (anonymous):

still have no idea what you mean?

OpenStudy (anonymous):

eq root 2 squared is 2

OpenStudy (anonymous):

\[\sqrt{number^2}=number\]

OpenStudy (anonymous):

True, but different to here...

OpenStudy (anonymous):

\[(\sqrt{number})^2=number\]

OpenStudy (anonymous):

\[\sqrt2^2 =2\]

OpenStudy (anonymous):

\[\sqrt{x-3}^2 =\]

OpenStudy (anonymous):

oh ok i got \[x=13 \] because \[\sqrt{13+3}=\sqrt{16}=\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

x-3 is the left hand side squared, agree?

OpenStudy (anonymous):

\[\sqrt{x+3}^{2}\]

OpenStudy (anonymous):

is that what you are asking

OpenStudy (anonymous):

What's that?

OpenStudy (anonymous):

Here is your question square root of x-3= 4

OpenStudy (anonymous):

yes that is the problem

OpenStudy (anonymous):

Do you agree that the square of the left hand side is x-3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x-3=16 x=19

OpenStudy (anonymous):

And the square of the right hand side is 16.

OpenStudy (anonymous):

mathballet is a genius!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

How did u do it?

OpenStudy (anonymous):

im even more confused

OpenStudy (anonymous):

\[\sqrt{x-3}=4\] \[ x-3= 4^{2}\]\[x-3=16\]

OpenStudy (anonymous):

I believe this is where we had got to before the interruption. x-3=16

OpenStudy (anonymous):

x=16+3 x=19

OpenStudy (anonymous):

duz it help

OpenStudy (anonymous):

SO mathballet used considerable math expertise in transferring the 3 to the other side.

OpenStudy (anonymous):

it is x+3 not x-3 in the sqrt

OpenStudy (anonymous):

crap i wrote it wrong

OpenStudy (anonymous):

but u posted sqrt x-3

OpenStudy (anonymous):

ok it is \[\sqrt{x+3}=4\] what is x

OpenStudy (anonymous):

That looks familiar, third time lucky?

OpenStudy (anonymous):

that is the problem sorry :D

OpenStudy (anonymous):

square it..

OpenStudy (anonymous):

=\[\sqrt{x+3}=16 \] yes?

OpenStudy (anonymous):

U have squared a square root, that get's rid of the square root.

OpenStudy (anonymous):

\[ x+3= 4^{2}\]

OpenStudy (anonymous):

x+3=16

OpenStudy (anonymous):

x=16-3 x=13

OpenStudy (anonymous):

so it is just \[x+3=16\] \[x=16-3=13 \] x=13

OpenStudy (anonymous):

yay i got it right thanks a lot

OpenStudy (anonymous):

\[\huge \checkmark\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Actually.. \[\sqrt{a^2} = |a|\]

OpenStudy (anonymous):

But it works out fine in this case.